OpenStudy (anonymous):

The crankshaft in a race car goes from rest to 3540rpm in 2.6s. What is the crankshaft's angular acceleration? How many revolutions does it make while reaching 3540rpm?

OpenStudy (3psilon):

use \[\omega = \omega _o + \alpha t\] \[\omega_o= 0 \] since it starts from rest. so \[\omega = \alpha t\] solve for alpha\[\alpha=\frac{ \omega }{ t }\] Now just plug in and solve. Also I'm assuming your answer should be in radians/sec? \[\frac{ 3540rev/m }{ 2.6\sec } * \frac{ 2\pi }{ 1 rev }*\frac{ 1 \min }{60 \sec } = 142.58 rad/s^{2}\]

OpenStudy (3psilon):

For the second part use the angular displacement formula \[\Delta \theta = \omega _o +\frac{ 1 }{ 2 } \alpha t^{2}\] and just plug in (note it starts from rest so omega initial = 0)

OpenStudy (anonymous):

Nice @3psilon

OpenStudy (3psilon):

Make sense @Ashley1593

OpenStudy (3psilon):

?

OpenStudy (anonymous):

yes thank you